I'm using The Complete Idiot's Guide to Pre-Algebra. I'm working on signed numbers and the answer to this one problem doesn't make sense.
-13 + 5= ? My answer was -8. The book says the answer is -7. Here is the explanation: This is the same as -(13-5)= -7. -13 is negative and 5 is positive, but 5 is less than 13, so you should expect the answer to be negative.

Found 2 solutions by Alan3354, MathTherapy:
Answer by Alan3354(69443)

-8 is correct.
Sometimes the book is wrong. Not often, but sometimes.

Answer by MathTherapy(10555)


I'm using The Complete Idiot's Guide to Pre-Algebra. I'm working on signed numbers and the answer to this one problem doesn't make sense.
-13 + 5= ? My answer was -8. The book says the answer is -7. Here is the explanation: This is the same as -(13-5)= -7. -13 is negative and 5 is positive, but 5 is less than 13, so you should expect the answer to be negative.

Subtracting 5 from 13 results in 8, after which the operation before the larger number (-) is kept.
How can 13 - 5 give 7? Are you sure you're looking at the right answer?